One-half the square of a number is equal to 16 more than twice the number. what is the number?
Answer by TimothyLamb(4379)

---
x^2/2 = 2x + 16
(1/2)x^2 - 2x - 16 = 0
0.5x^2 - 2x - 16 = 0
---
the above quadratic equation is in standard form, with a=0.5, b=-2, and c=-16
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
0.5 -2 -16
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
this quadratic has two real roots (the x-intercepts), which are:
x = 8
x = -4
---
answer:
two different numbers satisfy the problem statement, namely 8 and -4
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php